Circular No. 6771

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1997 V1 (LARSEN) As was suggested on IAUC 6770, further observations indicate that this

comet is of short period. Orbital elements (from observations extending to Nov. 9) on MPEC 1997-V23 (also MPC 30846) have P = 10.8 years.

COMET C/1997 V2 (SOHO) C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports that D. Lewis, University of Birmingham, has discovered another comet in both C3 and C2 coronagraphic data. The comet was very faint in C3, but since it was visible in C2, the magnitude was estimated at ∼ 6. There was likely a faint tail in some of the C2 images. Measurements by Lewis, D. A. Biesecker and St. Cyr and reduced by G. V. Williams are given in detail on MPEC 1997-V29; there is a particular abundance of measurements because the LASCO instruments were in high-cadence mode following a solar flare and proton event a week ago. The orbit computation on MPEC 1997-V29 by the undersigned is a full parabolic solution that shows the comet’s membership in the Kreutz sungrazing group.

1997 UT  α_{2000}  δ_{2000}
Nov. 8.578  14^{h}41.^{m}9  −19^{o}23′

SATELLITES OF URANUS Further CCD observations (again restricted here to one per observer

per night) are given below for S/1997 U 1 and S/1997 U 2, respectively. The likely general correctness of the S/1997 U 2 orbit on IAUC 6765 continues to be supported, but the S/1997 U 1 orbit is still pure speculation.

1997 UT α_{2000} δ_{2000} Observer
Oct. 30.84300 20^{h}29^{m}44.^{s}46 −19^{o}42′10′′.5 Fitzsimmons
Oct. 29.13938 20^{h}29^{m}12.^{s}34 −19^{o}35′31′′.6 Offutt
30.84300 20 29 18.35 −19 35 08.0 Fitzsimmons
Nov. 5.13278 20 29 40.83 −19 33 40.4 Offutt
8.13200 20 29 56.20 −19 32 42.8 Hergenrother
9.15096 20 30 01.76 −19 32 21.0 Offutt

A. Fitzsimmons, M. E. Fletcher and M. J. Irwin, assisted by P. Sorenson

and M. Asif (La Palma). 2.5-m Isaac Newton Telescope. W. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien. C. W. Hergenrother (Mt. Hopkins). 1.2-m f/8 reflector.

1997 November 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden